# BindWeave: Subject-Consistent Video Generation via Cross-Modal Integration

### v1

@@ 4.3 Qualitative results
To clearly demonstrate the effectiveness of our method, we present some typical subject-to-video scenarios in Figure 4 and Figure 5, including single-body-to-video, human-entity-to-video, single-object-to-video, and multi-entity-to-video. As shown in the left panel of Figure 4, commercial models such as Vidu, Pika, Kling, and Hailuo produce visually appealing videos but struggle with subject consistency. Among open-source methods, SkyReel-A2 is comparatively competitive on subject consistency, yet its overall visual aesthetics lag behind our BindWeave. VACE and Phantom similarly exhibit weak subject consistency. In the right panel of Figure 4, our approach achieves markedly better subject consistency, text alignment, and visual quality. As shown in the left panel of Figure 5, in single-object-to-video scenarios, commercial models such as Vidu and Pika still exhibit pronounced violations of physical and semantic plausibility—what we summarize as “commonsense violations” (e.g., a human walking with severely twisted legs). Kling achieves strong visual aesthetics but maintains poor subject consistency. SkyReels-A2 shows severe distortions and similarly weak subject consistency, and Phantom also struggles to preserve subject consistency. Among the baselines, VACE better maintains subject consistency but suffers from limited motion coherence and naturalness. In contrast, our BindWeave delivers strong subject consistency together with natural and coherent motion. Notably, under multi-object and complex-instruction settings as shown in the right panel of Figure 5, methods like Vidu and Pika often miss key cues (e.g., “hot oil”), Kling shows severe physical implausibility (e.g., fries leaking directly out of the basket), and MAGREF fails to preserve subject consistency; other baselines also omit crucial prompt details. In contrast, our results deliver fine-grained detail while maintaining strong subject consistency. We attribute this to BindWeave’s explicit cross-modal integration of the reference image and textual prompt via an MLLM, which jointly parses entities, attributes, and inter-object relations. As a result, BindWeave preserves subtle yet crucial details (e.g., “hot oil”) and constructs a unified, temporally consistent scene plan to guide coherent generation. This deep cross-modal integration reliably enforces key prompt elements and embeds basic physical commonsense for multi-entity interactions, thereby reducing implausible outcomes. More visualizations can be found in Appendix Section A.3, including additional qualitative examples and frame-wise comparisons.
@@ 4.4 Ablation study
We ablate our control-conditioning that concatenates MLLM- and T5-derived signals to guide a DiT during generation. We compare a T5-only baseline with our T5+Qwen2.5-VL variant, an MLLM-only setup (Qwen2.5-VL + DiT) was unstable and failed to converge within our training budget, so it is omitted from quantitative analysis. As shown in Table 2, T5+Qwen2.5-VL consistently outperforms T5-only across aesthetics, motion, naturalness, and text relevance. Qualitative comparisons in Figure 6 further corroborate these findings: when reference images exhibit scale mismatch, the T5-only baseline tends to produce unrealistic subject sizes (e.g., dog–bowl), and under complex instructions it often misparses action–object relations, whereas T5+Qwen2.5-VL remains well grounded and executes the intended interactions. We attribute these gains to complementary conditioning, the MLLM provides multimodal, identity- and relation-aware cues that disambiguate subjects and improve temporal coherence, while T5 offers precise linguistic grounding that stabilizes optimization. Their concatenation yields a richer and more reliable control signal for DiT.
@@ A.2 LLM Usage Disclosure
A.3 More subject-to-video quantitative results

### v2

@@ 4.3 Qualitative results
To clearly demonstrate the effectiveness of our method, we present some typical subject-to-video scenarios in Figure 4 and Figure 5, including single-body-to-video, human-entity-to-video, single-object-to-video, and multi-entity-to-video. As shown in the left panel of Figure 4, commercial models such as Vidu, Pika, Kling, and Hailuo produce visually appealing videos but struggle with subject consistency. Among open-source methods, SkyReel-A2 is comparatively competitive on subject consistency, yet its overall visual aesthetics lag behind our BindWeave. VACE and Phantom similarly exhibit weak subject consistency. In the right panel of Figure 4, our approach achieves markedly better subject consistency, text alignment, and visual quality. As shown in the left panel of Figure 5, in single-object-to-video scenarios, commercial models such as Vidu and Pika still exhibit pronounced violations of physical and semantic plausibility—what we summarize as “commonsense violations” (e.g., a human walking with severely twisted legs). Kling achieves strong visual aesthetics but maintains poor subject consistency. SkyReels-A2 shows severe distortions and similarly weak subject consistency, and Phantom also struggles to preserve subject consistency. Among the baselines, VACE better maintains subject consistency but suffers from limited motion coherence and naturalness. In contrast, our BindWeave delivers strong subject consistency together with natural and coherent motion. Notably, under multi-object and complex-instruction settings as shown in the right panel of Figure 5, methods like Vidu and Pika often miss key cues (e.g., “hot oil”), Kling shows severe physical implausibility (e.g., fries leaking directly out of the basket), and MAGREF fails to preserve subject consistency; other baselines also omit crucial prompt details. In contrast, our results deliver fine-grained detail while maintaining strong subject consistency. We attribute this to BindWeave’s explicit cross-modal integration of the reference image and textual prompt via an MLLM, which jointly parses entities, attributes, and inter-object relations. As a result, BindWeave preserves subtle yet crucial details (e.g., “hot oil”) and constructs a unified, temporally consistent scene plan to guide coherent generation. This deep cross-modal integration reliably enforces key prompt elements and embeds basic physical commonsense for multi-entity interactions, thereby reducing implausible outcomes. More visualizations can be found in Appendix Section A.5 and our supplementary materials, including additional qualitative examples and comparisons.
@@ 4.4 Ablation study
We ablate our control-conditioning that concatenates MLLM- and T5-derived signals to guide a DiT during generation. We compare a T5-only baseline with our T5+Qwen2.5-VL variant. The MLLM-only setup (Qwen2.5-VL + DiT) is omitted from this quantitative analysis because, as discussed in Section A.3, it proved to be unstable during training and failed to converge within our training budget. As shown in Table 2, T5+Qwen2.5-VL consistently outperforms T5-only across aesthetics, motion, naturalness, and text relevance. Qualitative comparisons in Figure 6 further corroborate these findings: when reference images exhibit scale mismatch, the T5-only baseline tends to produce unrealistic subject sizes (e.g., dog–bowl), and under complex instructions it often misparses action–object relations, whereas T5+Qwen2.5-VL remains well grounded and executes the intended interactions. We attribute these gains to complementary conditioning, the MLLM provides multimodal, identity- and relation-aware cues that disambiguate subjects and improve temporal coherence, while T5 offers precise linguistic grounding that stabilizes optimization. Their concatenation yields a richer and more reliable control signal for DiT. More cases are provided in Section A.4.
@@
4.5 User Study
We conducted a user study to evaluate our method, employing mean opinion scores (MOS) across four key criteria: subject consistency, prompt following, video quality, and motion quality. These dimensions provide a comprehensive assessment of the generated videos. We recruited 20 participants to perform anonymized comparisons across different methods. All generated samples are anonymized and randomly assigned to users, with each item scored on a 1–5 scale. The results are summarized in Table 3, and for clearer presentation we also visualize the data in Figure 7. The evaluation indicates that BindWeave achieves the best performance in subject consistency while maintaining leading results across all metrics.
Method
Subject Consistency $\uparrow$
Prompt Following $\uparrow$
Motion Quality $\uparrow$
Video Quality $\uparrow$
Total Score Average $\uparrow$
SkyReels 3.47 3.52 3.60 3.23 3.46
Vidu 3.40 3.55 3.63 3.50 3.52
Magref 3.25 3.65 3.65 3.58 3.53
Phantom 3.56 3.72 3.84 3.58 3.68
Kling 3.62 3.58 3.90 3.75 3.71
VACE 3.88 3.63 3.80 3.62 3.73
BindWeave (Ours) 3.94 3.66 3.75 3.70 3.76
Table 3: User study results comparing different methods. “Total Score” means the average score.
Figure 7: Visualization of user study scores across the evaluation criteria.
Figure 8: Evaluating copy–paste artifacts under conflict–coherence scenarios
4.6 Copy-paste problem
We evaluate whether conditioning on reference images induces pixel-level “copy-paste” artifacts by constructing conflict-coherence scenarios that deliberately mismatch the reference and the prompted outcome in Figure 8. Beyond contrasting facial expressions (e.g., smiling reference $\rightarrow$ painful, tearful video; painful reference $\rightarrow$ joyful smiling video), we also test outfit changes and diverse pose settings to more comprehensively validate our approach. Across these cases, the model follows the instructions rather than pasting pixels; it adapts facial musculature, attire, and pose to match the prompt while preserving identity, maintains smooth temporal transitions without stuck frames, and produces coherent motion instead of static overlays.
@@ 5 Conclusion
Acknowledgments
This work was partially supported by the National Key R&D Program of China (Grant No. 2024YFB3909902) and Youth Innovation Promotion Association of CAS.
@@ A.1 Appendix Overview
  • Section A.3: Justification for Model Design. In this section, we discuss our architectural choices and provide a clear, empirical justification for our final model design.
    Section A.4: Visual comparisons in complex multi-subject scenarios between T5-only and MLLM+T5. We present extended qualitative comparisons under complex multi-subject settings.
@@ A.2 LLM Usage Disclosure
A.3 Justification for Model Design
During the development phase, we conducted a series of pilot experiments to explore various architectural designs for integrating the Multi-modal Large Language Model (MLLM) and the Diffusion Transformer (DiT). These explorations served as an informal ablation study and directly informed our final design choice. Specifically, we investigated several alternatives to our proposed MLLM + MLP + T5 + DiT architecture, including:
  1. 1.
    MLLM + MLP + DiT: Using a simple Multi-Layer Perceptron (MLP) to project MLLM features into the DiT.
  2. 2.
    MLLM + Q-Former + DiT: Employing a Q-Former structure to refine the MLLM outputs before feeding them to the DiT.
Our key finding was that architectures relying solely on the MLLM for conditioning (i.e., those without the T5 text encoder) were unstable during training and failed to converge within our available training budget. These models struggled to effectively translate the nuanced visual identity information from the MLLM into the precise conditional inputs required by the DiT for high-fidelity video generation. In contrast, the inclusion of T5 provided a robust and stable text-conditioning backbone, which, when combined with the identity features from the MLLM, yielded the consistent and high-quality results presented in our paper. To provide concrete evidence of the convergence issues, we have included a training loss curve for the MLLM + MLP + DiT architecture (Figure 9). This curve clearly illustrates the training instability, showing that the model’s loss oscillates significantly and fails to reach convergence.
Figure 9: Training loss for the MLLM + MLP + DiT architecture, showing significant oscillation and failure to converge.
Figure 10: Qualitative comparison of T5-only vs. MLLM+T5.
A.4 Comparative results of T5-only versus MLLM+T5 (BindWeave) under complex scenarios
Figure 10 presents extensive challenging cases, including five‑image reference setups and complex multi‑subject interactions, with side‑by‑side comparisons between T5‑only and MLLM+T5. Across these multi-reference cases, the T5-only model frequently exhibits distortions and temporal jitter, leading to unstable videos. In contrast, MLLM+T5 reliably reasons about subject placement and overall layout, yielding coherent spatial arrangements, stable motion, and higher visual quality. In essence, T5 provides dense captions that help structure prompt interpretation and offer a stable descriptive prior. However, relying on T5 alone is insufficient for compositional spatial/temporal reasoning and multi-subject role disambiguation: T5-only often misses cross-entity relations, produces inconsistent event ordering, and may conflate subject roles—leading to layout instability and drift in videos. The MLLM supplies compositional reasoning and cross-entity constraint satisfaction. With MLLM+T5, the model correctly resolves ordered cross-entity relations and better maintains relative positions and action dependencies.
A.5 More subject-to-video quantitative results